Conduct a test to determine whether or not the population proportion of voters in favor of proposal A is greater than 50%. In a random sample of 200 voters, 140 said that they were in favor of this proposal. Compute the test statistic.
Please provide steps to arrive at the answer.

Answer by stanbon(75887)

Conduct a test to determine whether or not the population proportion of voters in favor of proposal A is greater than 50%. In a random sample of 200 voters, 140 said that they were in favor of this proposal. Compute the test statistic.
Please provide steps to arrive at the answer.
-----------
Ho: p <= 0.5
Ha: p > 0.5
----
test proportion: 140/200 = 0.7
------
test stat: z(0.7) = (0.7-0.5)/sqrt[0.5*0.5/200] = 5.66
-----
Conclusion: Since the test stat is in the reject interval,
reject Ho.
Conclusion: The test supports the conjecture that the proportion
is greater than 0.5.
--------------------------
Cheers,
Stan H.
==========================